OpenStudy (anonymous):

can some plz so steps on how to answer this. solve. (f)-2/3=5/6

OpenStudy (anonymous):

what is (f).@Wolverine32

OpenStudy (anonymous):

that's what we have to solve for /f/-2/3=5/6 this is absolute Value Equations and Inequalities

OpenStudy (anonymous):

Is it f(-2/3)= 5/6.

OpenStudy (anonymous):

\[|f|-2/3=5/6 this \implies |f|=5/6+2/3=3/2 this \implies f=\pm3/2\]

OpenStudy (anonymous):

@Wolverine32

OpenStudy (anonymous):

did you get it??

OpenStudy (anonymous):

no sorry had a little problem with open study